# HearHere: Mitigating Echo Chambers in News Consumption through an AI-based Web System

### 28 Feb 2024 v1

@@ Abstract.
Considerable efforts are currently underway to mitigate the negative impacts of echo chambers, such as increased susceptibility to fake news and resistance towards accepting scientific evidence. Prior research has presented the development of computer systems that support the consumption of news information from diverse political perspectives to mitigate the echo chamber effect. However, existing studies still lack the ability to effectively support the key processes of news information consumption and quantitatively identify a political stance towards the information. In this paper, we present HearHere, an AI-based web system designed to help users accommodate information and opinions from diverse perspectives. HearHere facilitates the key processes of news information consumption through two visualizations. Visualization 1 provides political news with quantitative political stance information, derived from our graph-based political classification model, and users can experience diverse perspectives (Hear). Visualization 2 allows users to express their opinions on specific political issues in a comment form and observe the position of their own opinions relative to pro-liberal and pro-conservative comments presented on a map interface (Here). Through a user study with 94 participants, we demonstrate the feasibility of HearHere in supporting the consumption of information from various perspectives. Our findings highlight the importance of providing political stance information and quantifying users’ political status as a means to mitigate political polarization. In addition, we propose design implications for system development, including the consideration of demographics such as political interest and providing users with initiatives.
@@ 2.2. AI to classify a political stance for mitigating echo chambers
Much research has been done for the purpose of developing AI models to classify political stances on news information. Feng et al. (Feng et al., 2021) presented KGAP, a political stance detection method that incorporates external domain knowledge to reflect the political context. Li et al. (Li and Goldwasser, 2019) suggested an AI model based on Graph Convolutional Networks for representing relational information, to capture the documents’ social context. Li and Goldwasser (Li and Goldwasser, 2021) proposed to MEAN (Multi-head Entity Aware Attention Network) which calculates the relationship between the entities in news and external information to determine polarization. Sun et al. (Sun et al., 2018) presented a neural model to fully employ various linguistic information to construct the document representation with an attention mechanism to protect the political stances of news articles.

### 29 Feb 2024 v2

@@ Abstract.
Considerable efforts are currently underway to mitigate the negative impacts of echo chambers, such as increased susceptibility to fake news and resistance towards accepting scientific evidence. Prior research has presented the development of computer systems that support the consumption of news information from diverse political perspectives to mitigate the echo chamber effect. However, existing studies still lack the ability to effectively support the key processes of news information consumption and quantitatively identify a political stance towards the information. In this paper, we present HearHere, an AI-based web system designed to help users accommodate information and opinions from diverse perspectives. HearHere facilitates the key processes of news information consumption through two visualizations. Visualization 1 provides political news with quantitative political stance information, derived from our graph-based political classification model, and users can experience diverse perspectives (Hear). Visualization 2 allows users to express their opinions on specific political issues in a comment form and observe the position of their own opinions relative to pro-liberal and pro-conservative comments presented on a map interface (Here). Through a user study with 94 participants, we demonstrate the feasibility of HearHere in supporting the consumption of information from various perspectives. Our findings highlight the importance of providing political stance information and quantifying users’ political status to mitigate political polarization. In addition, we propose design implications for system development, including the consideration of demographics such as political interest and providing users with initiatives.
@@ 2.2. AI to classify a political stance for mitigating echo chambers
Much research has been done for the purpose of developing AI models to classify political stances on news information. Feng et al. (Feng et al., 2021b) presented KGAP, a political stance detection method that incorporates external domain knowledge to reflect the political context. Li et al. (Li and Goldwasser, 2019) suggested an AI model based on Graph Convolutional Networks for representing relational information, to capture the documents’ social context. Li and Goldwasser (Li and Goldwasser, 2021) proposed to MEAN (Multi-head Entity Aware Attention Network) which calculates the relationship between the entities in news and external information to determine polarization. Sun et al. (Sun et al., 2018b) presented a neural model to fully employ various linguistic information to construct the document representation with an attention mechanism to protect the political stances of news articles.
@@ Acknowledgements.
Appendix A Appendix
In this appendix, we provide a detailed description of the user study for modeling (Appendix A.1), development of our model (Appendix A.2), the process of constructing the data (Appendix A.3), and the additional experimental results on the reliability of political stance prediction experiments (Appendix A.4).
A.1. User Study
In this section, we present details of our user study which aims to investigate the key factors considered by real-life users when evaluating the political stance expressed in news articles. We first conducted a user study in order to investigate what factors real-world users consider and how important the factors are in determining the political stances of news articles. We provided six articles with different political stances and carefully-chosen factors to 136 respondents (69 male, 67 female) via Amazon Mechanical Turk. We asked them to respond with how important each factor is in their decision with a scale [1: not at all, 5: very much]. The ”context” of a new article is the most important factor in deciding its political stance, followed by keywords, person, tone, and frequently used words. Based on this result, we observe that it is crucial (Observation 1) to learn the relationships among words/sentences to capture the context and tone of a news article, which is implicitly reflected in the article, and (Observation 2) to understand the interpretation and sentiment to real-world entities (e.g., keyword and person), which explicitly appear in a news article.
User study setup.
A total of 136 participants were recruited for our user study through Amazon Mechanical Turk  [^3] https://www.mturk.com/. Our selection process took into account diverse criteria including gender (male/female), age (ranging from between 20 to above 50), education level (ranging from high school or lower to university graduation and beyond), ethnicity (including but not limited to Caucasian, African American, Asian, Hispanic/Latino, and others), and political stance (ranging from very liberal to somewhat liberal, neutral, somewhat conservative, and very conservative).
To investigate the influence of political stances, we carefully selected six news articles that represented different political positions. Additionally, we considered thirteen political-related factors that have been previously examined in  (Lin et al., 2006; Gottopati et al., 2013; Gentzkow and Shapiro, 2010; Kuttschreuter et al., 2011; Irvine et al., 1992), These factors encompassed various aspects such as context and tone, keywords, names of individuals, topics and issues addressed, article titles, images accompanying the news articles, usage of slang words, and social and religious influences. Moreover, we focused on three news topics— Health, Environment, and Tax— as they are highly interconnected with political stances.
User study protocol.
Subsequently, we proceeded with the following steps: (1) We provided the participants with six news articles, including the title, body, and image, along with the thirteen selected factors. (2) We asked the participants to indicate the level of significance attributed to each factor in their decision-making process, using a Likert-scale ranging from 1 (not at all significant) to 5 (very significant). (3) We evaluated the extent to which each factor helped identify the political stance by calculating the average scores assigned by the participants.
Result and analysis.
The study results revealed that the most significant factor in determining the political stance is the ”context” of the article, followed by keywords, individuals mentioned, tone, and frequently used words. These results emphasize the importance of two key aspects: (1) understanding the interrelationships between words and sentences to capture the context and tone that are implicitly conveyed in the news article, and (2) identifying the interpretation and sentiment associated with real-world entities, such as keywords and individuals, which are explicitly present in the article.
A.2. Model development
We developed an AI model to predict the political stance of news articles by reflecting the results of our user study on building the model. The model consists of two key components: (1) hierarchical attention networks to learn the relationships among words/sentences in a news article with a three-level hierarchy (i.e., word-level, sentence-level, and title-level) and (2) knowledge encoding to incorporate both common and political knowledge for real-world entities, necessary for deeply understanding a news article, into the process of predicting the political stance of a news article. Regarding political knowledge, the interpretation and sentiment of even the same entities can also be different depending on their political stance. To address the subtle but important difference, we construct two knowledge graphs (KG) with different political stances, KG-lib and KG-con, and design knowledge encoding to learn to fuse the information extracted from the two political knowledge graphs. We trained the model for 50 epochs using a large dataset of news articles with five political stance labels (left, lean left, center, lean right, and right).
A.3. Data Construction
In this section, we describe the process to construct the datasets used in our study, namely: (1) the large-scale political news dataset, AllSides-L, and (2) two political knowledge graphs, KG-lib and KG-con.
AllSides-L.
We complied a comprehensive political news dataset, AllSides-L. This dataset consists of 719,256 articles categorized into five classes: left, lean left, center, lean right, and right. The articles were collected from Allsides.com  [^4] https://www.allsides.com/, an American website dedicated to mitigating the adverse effects of media bias and misinformation. Allsides.com offers a wide range of political news articles that cover diverse political stances. The political classification of each news article is determined based on its associated news outlet, such as CNN or Fox. Allsides.com employs a meticulous three-step process to assign political labels: (1) domain experts label each news outlet, (2) user studies involving individuals with varying political perspectives, and (3) majority voting by individuals not involved in the user studies. Due to this rigorous labeling process, the political stances assigned by Allsides.com are generally considered as the ground truth in previous studies (Feng et al., 2021a; Zhang et al., 2022). To ensure the reliability of AllSides-L, we only included news articles from outlets that received high scores (7 out of 8 or higher) in the majority voting process.
KG-lib and KG-con.
We constructed two distinct political knowledge graphs (KGs), KG-lib and KG-con, following a three-step process: (1) data collection, (2) entity/relation extraction, and (3) data cleansing. Initially, we collected a total of 496,071 posts, with 219,915 posts originating from the U.S. liberal community and 276,156 posts from the U.S. conservative community. Since these raw posts may contain numerous entities unrelated to politics, our objective was to extract political entities and their corresponding relationships from the posts. To achieve this, we employed a state-of-the-art Named Entity Recognition (NER) method (Vychegzhanin and Kotelnikov, 2019) to identify 18 political-related entities and their associated relations. Each data point was then represented as a triplet: ¡head entity, relation, tail entity¿. To ensure the reliability of the political knowledge graphs, we manually removed any extraneous noise from the extracted triplets. As a result, we constructed two political knowledge graphs: KG-lib consisting of 5,581 entities and 29,967 relations, and KG-con comprising 6,316 entities and 33,207 relations.
Table 7. Comparison of the model accuracy on three real-world datasets (The bold font indicates the best results).
Method Dataset
SemEval AllSides-S AllSides-L
Word2Vec (Mikolov et al., 2013) 0.7027 0.4858 0.4851
GloVe (Pennington et al., 2014) 0.8071 0.7101 0.6354
ELMo (Peters et al., 2018) 0.8678 0.8197 0.7483
BERT (Kenton and Toutanova, 2019) 0.8692 0.8246 0.7812
RoBERTa (Liu et al., 2019) 0.8708 0.8535 0.8222
KGAP (Feng et al., 2021a) 0.8956 0.8602 N/A
KCD (Zhang et al., 2022) 0.9087 0.8738 N/A
Ours-RotatE 0.9426 0.9151 0.8584
Ours-HAKE 0.9395 0.9216 0.8563
Ours-ModE 0.9521 0.9256 0.8617
A.4. Accuracy of our AI model
In this section, we aim to evaluate our political stance prediction model used in our empirical evaluation, with a specific focus on the accuracy of our model.
Evaluation protocol.
To assess the accuracy of our AI model, we adopted the following evaluation protocol. We compared our model with seven baseline methods using three datasets: SemEval, AllSides-S, and AllSides-L: five text-based methods (Mikolov et al., 2013; Pennington et al., 2014; Peters et al., 2018; Kenton and Toutanova, 2019; Liu et al., 2019) and two knowledge-based political stance prediction methods (Feng et al., 2021a; Zhang et al., 2022). Word2Vec (Mikolov et al., 2013), GloVe (Pennington et al., 2014), and ELMo (Peters et al., 2018) are general language models that aim to capture context from text. We also use pre-trained BERT (Kenton and Toutanova, 2019) and RoBERTa (Liu et al., 2019) by fine-tuning on training datasets. KGAP (Feng et al., 2021a) is a knowledge-aware approach that leverages a political knowledge graph with graph neural networks. KCD (Zhang et al., 2022), the state-of-the-art model, considers knowledge walks from a political knowledge graph and textual cues in political stance prediction. To ensure consistency, we (1) implemented the five baseline language models using their available source code, incorporating a softmax layer for the final political stance prediction. Subsequently, we (2) applied the baseline methods to the SemEval, AllSides-S, and AllSides-L datasets, (3) calculated their political stance prediction accuracies, and (4) reported the obtained results. For the two knowledge-based methods (i.e., KGAP (Feng et al., 2021a) and KCD (Zhang et al., 2022)), however, we were unable to obtain their results due to the unavailability of certain parts of the source code. Specifically, the source code for generating cue embeddings in KCD and the source code for generating knowledge embeddings in KGAP were not provided, respectively. Thus, we used the results of KGAP and KCD on the SemEval and AllSides-S datasets, reported in (Zhang et al., 2022).  [^5] KGAP: https://github.com/BunsenFeng/news_stance_detection, KCD: https://github.com/Wenqian-Zhang/KCD
Results and analysis.
We train our AI model on the three real-world political news article datasets for 50 epochs with varying knowledge embedding methods (RotatE (Sun et al., 2018a), HAKE (Zhang et al., 2020), and ModE (Zhang et al., 2020)). For the SemEval and AllSides-L datasets, we apply k-fold cross validations (10-fold for SemEval and 3-fold for AllSides-S) and then report the averaged accuracy of k-fold cross validations. For AllSides-L, we train the model on the training set (647k articles) and measure the model accuracy using the validation set (71.9k articles), which is not used in the training process. Tables 7 presents the results. Our AI model consistently outperforms all baseline methods in terms of the model accuracy, regardless of knowledge embedding methods. Specifically, Our model improves the state-of-the-art method, KCD (Zhang et al., 2022) by 4.77% and 5.92% in SemEval and AllSides-S datasets, respectively. These improvements over KCD are significant, given that KCD has already achieved quite high accuracies in those datasets. We also evaluate our model on a large-scale dataset (AllSides-L) which is 48×48$\times$48 × larger and has more classes (i.e., more difficult to predict) than AllSides-S. our model still significantly outperforms all baseline methods although the accuracy of our model decreases to some extent in AllSides-L, compared with the other datasets. In addition, we have conducted the t-tests with a 95% confidence level and verified that the improvement of our model over all baseline methods are statistically significant (i.e., the p-values are below 0.05). As a result, considering the consistent superiority of our model to overall baseline methods, our experimental results convincingly demonstrate the superiority of our model in comparison to existing political stance prediction methods.